OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable z is eliminated in the first and second equations, then the second and third equations. x + y – 3z = –8 2x + 2y + z = 12 3x + y – z = –2

OpenStudy (anonymous):

@johnweldon1993 can you help??

OpenStudy (anonymous):

A x + y = –8 3x + y = –2 B x + y = –8 5x + 3y = 14 C 7x + 7y = 28 3x + y = –2 D 7x + 7y = 28 5x + 3y = 10

OpenStudy (johnweldon1993):

hmm okay....well in order to eliminate it in the first and second equations....we need to multiply that 2nd equation by 3 2x + 2y + z = 12 becomes 6x + 6y + 3z = 36 so now we have (the first 2 are) x + y – 3z = –8 6x + 6y + 3z = 36 When we combine them we get 7x + 7y = 28 So now to eliminate the z in the third equation from the second equation we just combine them 2x + 2y + z = 12 3x + y – z = –2 After combining (and by that I mean adding them together) we have 5x + 3y = 10 so those 2 again are 7x + 7y = 28 5x + 3y = 10 Look familiar again? :)

OpenStudy (anonymous):

yess it does thankyou so much! :)

OpenStudy (johnweldon1993):

No problem! Time to go eat just like primeralph lol :)

OpenStudy (anonymous):

lol okaay! idk what im gonna do now...